OpenStudy (anonymous):

Find the magnitude and direction of the resultant of two forces of 250 pounds and 45 pounds at angles of 25° and 250° with the x-axis, respectively.

OpenStudy (anonymous):

 so since R^2=(f1)^2+(f2)^2+2(f1.f2)cosAlpha so R^2=(250)^2+(45)^2+2(11250)cos225 therefore: R^2=48615.09 so R=220.48 Pounds